OpenStudy (dottiecat):

Hadar went shopping for supplies for her dog. She spent $14 for dog food, $5 for a collar, and $11 for a leash. What percent of the amount that Hadar spent on the supplies was the price of the leash? Round to the nearest percent, if necessary.

OpenStudy (sbuck98):

14+5= 19+11=30 30*11%=3.30

OpenStudy (dottiecat):

Thank you very much, but I would've liked if you helped me. not give me the answer :(